# Artifact-free Sound Quality in DNN-based Closed-loop Systems for Audio Processing

### v2

@@ Abstract
Recent advances in deep neural networks (DNNs) have significantly improved various audio processing applications, including speech enhancement, synthesis, and hearing aid algorithms. DNN-based closed-loop systems have gained popularity in these applications due to their robust performance and ability to adapt to diverse conditions. Despite their effectiveness, current DNN-based closed-loop systems often suffer from sound quality degradation caused by artifacts introduced by suboptimal sampling methods. To address this challenge, we introduce dCoNNear, a novel DNN architecture designed for seamless integration into closed-loop frameworks. This architecture specifically aims to prevent the generation of spurious artifacts. We demonstrate the effectiveness of dCoNNear through a proof-of-principle example within a closed-loop framework that employs biophysically realistic models of auditory processing for both normal and hearing-impaired profiles to design personalized hearing aid algorithms. Our results show that dCoNNear not only accurately simulates all processing stages of existing non-DNN biophysical models but also eliminates audible artifacts, thereby enhancing the sound quality of the resulting hearing aid algorithms. This study presents a novel, artifact-free closed-loop framework that improves the sound quality of audio processing systems, offering a promising solution for high-fidelity applications in audio and hearing technologies.
@@
deep learning, closed loop, audio processing, artifacts, sound quality.
@@ I Introduction
Recent advances in deep neural networks (DNN) have demonstrated remarkable success in various audio processing applications, such as speech synthesis [1, 2, 3], speech recognition [4], speech enhancement [5], hearing aid algorithms [6], and others. DNN-based closed-loop frameworks have become increasingly prevalent across various audio applications, primarily due to their notable performance improvements and high flexibility in adapting to diverse conditions. The generalized framework illustrated in Fig. 1 integrates an audio processor with a condition module that can be customized for different tasks. These include using generative adversarial networks (GAN) for audio synthesis [1, 2, 3], employing DNN-based loss functions for speech enhancement [7, 8, 9, 10], and implementing biophysically inspired closed-loop frameworks for designing individualized hearing aid algorithms [6].
@@
However, current closed-loop frameworks are corrupted by the non-ideal downsampling and upsampling processes between the NN layers. The problematic operations can generate undesired artifacts, which degrade the resulting audio quality [11]. Specifically, typical downsampling methods, such as strided convolutions and average pooling, often lack low-pass filtering, leading to aliasing artifacts in low-frequency bands [12, 3]. Aliasing artifacts resulting from non-ideal downsampling methods degrade the harmonic components in the speech synthesis, leading to a noticeable decline in perceptual quality [13]. On the other hand, upsampling methods, such as transposed convolutions and pixel convolutions, tend to generate tonal artifacts that introduce noise in the high-frequency bands [11, 3]. Additionally, upsampling operations can produce imaging artifacts, where low frequencies are mirrored in high-frequency bands due to spectral replicas [3, 14]. For instance, [1] applied GANs to unsupervised audio generation. However, the use of transposed convolutions introduced pitched noise in the synthesized audio samples. Furthermore, [6] proposed a bio-inspired DNN-based closed-loop framework for designing hearing-aid (HA) algorithms. The autoencoder-based model (CoNNear) that is used within the closed-loop system comprises a differentiable description of the cochlea, inner-hair-cell (IHC), and auditory-nerve fiber (ANF) processing stages. However, the network produces audible artifacts that compromise the sound quality of the HA model. In these autoencoder-based auditory models (CoNNearcochlear, CoNNearIHC and CoNNearANF), the undesired artifacts originate from the transposed convolutions in the decoder.
@@
To eliminate aliasing during upsampling and downsampling, Gaussian blur and low-pass filters can be used, as adopted in Alias-CNN [12] and stylegan3 [15], respectively. However, these methods resulted in reduced performance in speech synthesis tasks [14]. [14] modified the transposed convolution layer to blend high-frequency features from the input with low-pass filtered features at each upsampling stage. It showed potential in maintaining high performance while reducing aliasing artifacts. Alternative upsampling methods like linear interpolation have been proposed to eliminate the tonal artifacts [16, 17]. The linear interpolation can solve the problem of tonal artifacts, but still leaves imaging artifacts in the high-frequency bands [14].
@@
We propose an artifact-free architecture (dCoNNear) designed to integrate the closed-loop systems for audio applications. We will illustrate the approach for hearing aid algorithms [6] with high sound quality. The dCoNNear is inspired by temporal convolutional networks (TCN) [18] and deep feedforward sequential memory networks (DFSMN) [19]. It comprises a sequence of stacked memory blocks. For each memory block, depthwise dilated 1-D convolutions are employed to model the long-term dependencies of auditory and audio processing (e.g. cochlear impulse response and neuronal adaptation) while avoiding the need for downsampling and upsampling in prior CoNNear-based systems. The dCoNNear architecture is applied to all auditory processing stages in the closed-loop system, including cochlear, IHC, and ANF, as well as the sound processor to train a hearing-aid signal processing algorithm (see Fig. 2). We show that dCoNNear accurately simulates all processing stages of non-DNN-based SOTA biophysical auditory processing without introducing spurious and audible artifacts in the resulting closed-loop trained audio application. We further extend the approach to a generalised artifact-free method for DNN-based closed-loop audio applications with high sound quality.
@@
The rest of the paper is organized as follows: Section II provides an in-depth characterization of the artifacts. Section III introduces the proposed artifact-free dCoNNear-based closed-loop system. Section IV details the experimental procedures, while Section V outlines the evaluation methods. Sections VI and VII present the results and discussions, respectively. Finally, Section VIII concludes the study.
@@ II Characteristics of upsampling Artifacts
The upsampling artifacts are well documented, [11] has demonstrated that transposed convolutions and subpixel convolutions introduce tonal artifacts, while nearest neighbor interpolation is prone to generating filtering artifacts. Additionally, during the upsampling process, the spectrum is cyclically repeated at the sampling rate, causing low-frequency components to be mirrored into high-frequency bands. This phenomenon leads to the replication of artifacts introduced by problematic upsampling methods, and the unintended presence of low-frequency components in high-frequency bands, commonly referred to as imaging artifacts. However, [11] analyzed only the artifacts after weight initialization, without providing a systematic analysis of the impact of the learning process on upsampling artifacts. In this section, we systematically characterize the artifacts generated by transposed convolutions, subpixel convolutions, and nearest neighbor interpolation when training the network for modeling auditory processing. We then studied artifacts associated with transposed convolution within the context of the closed-loop, CNN-based hearing aid algorithms [6].
@@
To characterize the upsampling artifacts when employing neural networks to emulate the auditory model, we investigated the artifacts of the autoencoder-based CoNNearcochlear [20] with 4 encoders and decoders, herein referred to as prior CoNNear, between different upsampling methods. CoNNearcochlear is a neural network representation of a non-linear transmission-line (TL) model that faithfully simulates the basilar-membrane (BM) displacement of human cochlear processing [21]. We examined three upsampling strategies in the decoders: transposed convolutions, subpixel convolutions, and nearest-neighbor interpolation. As depicted in Fig. 3, transposed and subpixel convolutions showed the peaks that were absent in the target model while the nearest-neighbor interpolation showed no obvious extra peaks in the step responses. To mitigate aliasing artifacts introduced by the strided convolution layers, which have a step size of 2, we applied a low-pass filter with a normalized cutoff frequency of 0.5 before each downsampling layer. All upsampling methods, however, showed additional spectral peaks in response to the 1-kHz tone. Previous studies have identified periodic tonal artifacts in transposed and subpixel convolutions due to problematic upsampling operators and spectral replicas after weight initialization [11]. In our findings, these artifacts persisted even after training. While nearest-neighbor interpolation is known to avoid tonal artifacts, it failed to effectively suppress imaging artifacts, resulting in noticeable peaks in the 1-kHz tone responses after training. These results suggest that the training process did not sufficiently resolve the artifacts caused by architectural limitations and spectral replicas, resulting in undesirable distortions in the model’s output.
@@
The artifacts associated with the autoencoder-based CoNNear models were systematically examined across each auditory processing stage depicted in Fig. 2, which include CoNNearcochlear, CoNNearIHC and CoNNearANF stages as part of the closed-loop system. These artifacts are identified and quantified using the deterministic auditory computational models [22], which served as targets during prior CoNNear training. Frequency responses for a 1-kHz tonal input across these stages are illustrated in Fig. 4(a-c). Fig. 4d shows the output of the HA model trained with the prior CoNNear framework to compensate for a high-frequency sloping hearing loss [6]. The prior CoNNear-based models exhibit spectral peaks in the magnitude spectrum of the cochlear, IHC, and ANF responses that are not present in the target model. During the training of the HA model, which minimizes the difference between NH and HI AN responses, the HA models incorporated tonal artifacts as shown in Fig. 4d. These artifacts degrade the resulting audio quality and should be systematically excluded from the closed-loop framework.
@@
Figure 2: Diagram of the closed-loop framework for designing individualized hearing aid algorithms.
@@ III Artifact-free dCoNNear-based closed-loop system
The artifact-free dCoNNear-based closed-loop framework for individualized DNN-based hearing aid (HA) model training is illustrated in Fig. 2. This framework features two pathways: one for the AN response ($r_{f}$) of a normal hearing (NH) system, and the other for the response ($\hat{r_{f}}$) of a hearing-impaired (HI) system. Each pathway includes three primary components corresponding to distinct auditory processing stages, the cochlea, inner hair cells (IHC), and auditory nerve fibers (ANF). The NH AN response ($r_{f}$) is generated by combining three types of fibers (high-spontaneous rate (HSR), medium-spontaneous rate (MSR), and low-spontaneous rate (LSR)). These fibers are weighted by $H_{NH}$, $M_{NH}$, and $L_{NH}$, reflecting the typical innervation of each IHC in a healthy auditory system. To simulate hearing-impaired profiles, we adjust these weights to $H_{HI}$, $M_{HI}$, and $L_{HI}$ to reflect the effects of cochlear synaptopathy (CS). The resulting AN responses ($r_{f}$) and ($\hat{r_{f}}$) provide biophysically realistic time-frequency representations of sound (neurograms), simulated at different cochlear locations. These neurograms reflect instantaneous firing rates across cochlear channels with center frequencies (CFs) ranging from 112 Hz to 12 kHz [23]. To tailor the hearing-impaired periphery to an individual’s sensorineural hearing loss (SNHL) profile, adjustments are made using audiometry to simulate the outer-hair-cell (OHC) damage in the cochlea or using auditory evoked potentials [24] to estimate CS. This involves introducing frequency-dependent OHC loss in CoNNearcochlear and/or CS in the CoNNearANF model. The DNN-based HA models are subsequently trained by minimizing a predefined loss function between the simulated NH and HI responses [6]. All auditory elements and HA models within this framework are built upon the dCoNNear topology.
@@
Inspired by temporal convolutional networks (TCN) [25] and deep feedforward sequential memory networks (DFSMN) [19], we introduce the dCoNNear architecture, as shown in Fig. 5. The dCoNNear incorporates a series of stacked memory blocks that function similarly to a high-order finite impulse response (FIR) filter [19], enabling the network to capture extensive dependencies in sequential signals. The model employs M memory blocks with dilation factors $2^{m-1}$, repeated R times, resulting in a total of N = M * R blocks. Residual connections between memory blocks facilitate learning by addressing the vanishing gradient problem. Skip connections aggregate the outputs from each memory block through weighted summation, and these aggregated outputs are then processed by a non-linear activation function followed by a convolutional layer to produce the final model outputs [25].
@@
The memory block (as shown in Fig. 5b) processes inputs from the previous block through pointwise convolution, followed by dilated depthwise convolution layers, Dh-conv and Df-conv, which capture historical and future context information. Depthwise convolution applies a single filter per input channel, reducing the number of parameters and helping capture long-term dependencies while maintaining a compact model size. The outputs from the pointwise and depthwise convolution layers are combined, passed through an activation function, and directed to subsequent blocks and skip connection paths. The learnable FIR-like memory block [19] is formulated as:

### v3

# dCoNNear: An Artifact-Free Neural Network Architecture for Closed-loop Audio Signal Processing

@@ Abstract
Recent advances in deep neural networks (DNNs) have significantly improved various audio processing applications, including speech enhancement, synthesis, and hearing-aid algorithms. DNN-based closed-loop systems have gained popularity in these applications due to their robust performance and ability to adapt to diverse conditions. Despite their effectiveness, current DNN-based closed-loop systems often suffer from sound quality degradation caused by artifacts introduced by suboptimal sampling methods. To address this challenge, we introduce dCoNNear, a novel DNN architecture designed for seamless integration into closed-loop frameworks. This architecture specifically aims to prevent the generation of spurious artifacts-most notably tonal and aliasing artifacts arising from non-ideal sampling layers. We demonstrate the effectiveness of dCoNNear through a proof-of-principle example within a closed-loop framework that employs biophysically realistic models of auditory processing for both normal and hearing-impaired profiles to design personalized hearing-aid algorithms. We further validate the broader applicability and artifact-free performance of dCoNNear through speech-enhancement experiments, confirming its ability to improve perceptual sound quality without introducing architecture-induced artifacts. Our results show that dCoNNear not only accurately simulates all processing stages of existing non-DNN biophysical models but also significantly improves sound quality by eliminating audible artifacts in both hearing-aid and speech-enhancement applications. This study offers a robust, perceptually transparent closed-loop processing framework for high-fidelity audio applications.
@@
deep learning, closed loop, audio signal processing, artifacts, sound quality.
@@ I Introduction
Recent advances in deep neural networks (DNN) have demonstrated remarkable success in various audio processing applications, such as speech synthesis [donahue2018adversarial, kong2020hifi, bak2023avocodo], speech recognition [abdel2014convolutional], speech enhancement [park2017ConvSE], hearing-aid algorithms [drakopoulos2023neural, drakopoulos2023dnn, drakopoulos2022differentiable], and others. DNN-based closed-loop frameworks have become increasingly prevalent across various audio applications, primarily due to their notable performance improvements and high flexibility in adapting to diverse conditions. The generalized framework illustrated in Fig. 1 integrates an audio processor with a condition module that can be customized for different tasks. These include using generative adversarial networks (GAN) for audio synthesis [donahue2018adversarial, kong2020hifi, bak2023avocodo], employing DNN-based loss functions for speech enhancement [germain2019speech, SEQualityNet, xu2021deep, fu2022metricgan], and implementing biophysically inspired closed-loop frameworks for designing individualized hearing-aid algorithms [drakopoulos2023neural, drakopoulos2023dnn, drakopoulos2022differentiable].
@@
However, current closed-loop frameworks are corrupted by the non-ideal downsampling and upsampling processes between the NN layers. The problematic operations can generate undesired artifacts, which degrade the resulting audio quality [pons2021upsampling]. Specifically, typical downsampling methods, such as strided convolutions and average pooling, often lack low-pass filtering, leading to aliasing artifacts in low-frequency bands [zhang2019making, bak2023avocodo]. Aliasing artifacts resulting from non-ideal downsampling methods degrade the harmonic components in the speech synthesis, leading to a noticeable decline in perceptual quality [zaidi2022daft]. On the other hand, upsampling methods, such as transposed convolutions and pixel convolutions, tend to generate tonal artifacts that introduce noise in the high-frequency bands [pons2021upsampling, bak2023avocodo]. Additionally, upsampling operations can produce imaging artifacts, where low frequencies are mirrored in high-frequency bands due to spectral replicas [bak2023avocodo, shang2023analysis]. For instance, [donahue2018adversarial] applied GANs to unsupervised audio generation. However, the use of transposed convolutions introduced pitched noise in the synthesized audio samples. Furthermore, [drakopoulos2023neural] proposed a bio-inspired DNN-based closed-loop framework for designing hearing-aid (HA) algorithms. The autoencoder-based model (CoNNear) that is used within the closed-loop system comprises a differentiable description of the cochlea, inner-hair-cell (IHC), and auditory-nerve fiber (ANF) processing stages. However, the network produces audible artifacts that compromise the sound quality of the HA model. In these autoencoder-based auditory models (CoNNearcochlear, CoNNearIHC and CoNNearANF), the undesired artifacts originate from the transposed convolutions in the decoder.
@@
To eliminate aliasing during upsampling and downsampling, Gaussian blur and low-pass filters can be used, as adopted in Alias-CNN [zhang2019making] and stylegan3 [karras2021alias], respectively. However, these methods resulted in reduced performance in speech synthesis tasks [shang2023analysis]. [shang2023analysis] modified the transposed convolution layer to blend high-frequency features from the input with low-pass filtered features at each upsampling stage. It showed potential in maintaining high performance while reducing aliasing artifacts. Alternative upsampling methods like linear interpolation have been proposed to eliminate the tonal artifacts [odena2016deconvolution, stoller2018waveUNet]. The linear interpolation can solve the problem of tonal artifacts, but still leaves imaging artifacts in the high-frequency bands [shang2023analysis].
@@
We propose dCoNNear, a novel architecture aimed at minimizing tonal and aliasing artifacts that commonly arise from suboptimal downsampling and upsampling operations in closed-loop DNN systems. We illustrate the approach for hearing-aid algorithms [drakopoulos2023neural] with high sound quality. Beyond hearing-aid applications, we demonstrate the effectiveness of dCoNNear in a closed-loop speech enhancement task, confirming its broader applicability to diverse wave-to-wave audio processing tasks. The dCoNNear is inspired by temporal convolutional networks (TCN) [lea2016temporal] and deep feedforward sequential memory networks (DFSMN) [zhang2018DFSMN]. It comprises a sequence of stacked memory blocks. For each memory block, depthwise dilated 1-D convolutions are employed to model the long-term dependencies of auditory and audio processing (e.g. cochlear impulse response and neuronal adaptation). This design eliminates the need for downsampling and upsampling, addressing limitations in prior CoNNear-based systems. The dCoNNear architecture is applied across all auditory processing stages in the closed-loop system, including cochlear, IHC, and ANF, as well as the sound processor for both hearing-aid and speech-enhancement algorithms. We show that dCoNNear accurately simulates all processing stages of non-DNN-based SOTA biophysical auditory processing, without introducing spurious and audible artifacts in the resulting closed-loop trained audio applications.
@@
The paper is organized as follows: Section II provides an in-depth characterization of the artifacts. Section III introduces the proposed artifact-free dCoNNear-based closed-loop system. Section IV details the experimental procedures, while Section V outlines the evaluation methods. Section VI presents the results for the hearing-aid task, while Section VII demonstrates the speech-enhancement application. Section VIII discusses the findings, and Section IX presents a conclusion.
@@ II Characteristics of upsampling Artifacts
The upsampling artifacts are well documented, [pons2021upsampling] has demonstrated that transposed convolutions and subpixel convolutions introduce tonal artifacts, while nearest neighbor interpolation is prone to generating filtering artifacts. Additionally, during the upsampling process, the spectrum is cyclically repeated at the sampling rate, causing low-frequency components to be mirrored into high-frequency bands. This phenomenon leads to the replication of artifacts introduced by problematic upsampling methods, and the unintended presence of low-frequency components in high-frequency bands, commonly referred to as imaging artifacts. However, [pons2021upsampling] analyzed only the artifacts after weight initialization, without providing a systematic analysis of the impact of the learning process on upsampling artifacts. In this section, we systematically characterize the artifacts generated by transposed convolutions, subpixel convolutions, and nearest neighbor interpolation when training the network for modeling auditory processing. We then studied artifacts associated with transposed convolution within the context of closed-loop, CNN-based hearing-aid algorithms [drakopoulos2023neural].
@@
To characterize the upsampling artifacts when employing neural networks to emulate the auditory model, we investigated the artifacts of the autoencoder-based CoNNearcochlear [baby2021convolutional] with 4 encoders and decoders, herein referred to as prior CoNNear, between different upsampling methods. CoNNearcochlear is a neural network representation of a non-linear transmission-line (TL) model that faithfully simulates the basilar-membrane (BM) displacement of human cochlear processing [verhulst2012TL]. We examined three upsampling strategies in the decoders: transposed convolutions, subpixel convolutions, and nearest-neighbor interpolation. As depicted in Fig. 3, transposed and subpixel convolutions showed the peaks that were absent in the target model while the nearest-neighbor interpolation showed no obvious extra peaks in the step responses. To mitigate aliasing artifacts introduced by the strided convolution layers, which have a step size of 2, we applied a low-pass filter with a normalized cutoff frequency of 0.5 before each downsampling layer. All upsampling methods showed additional spectral peaks in response to the 1-kHz tone. Previous studies have identified periodic tonal artifacts in transposed and subpixel convolutions due to problematic upsampling operators and spectral replicas after weight initialization [pons2021upsampling]. In our findings, these artifacts persisted even after training. While nearest-neighbor interpolation is known to avoid tonal artifacts, it failed to effectively suppress imaging artifacts, resulting in noticeable peaks in the 1-kHz tone responses after training. These results suggest that the training process did not sufficiently resolve the artifacts caused by architectural limitations and spectral replicas, resulting in undesirable distortions in the model’s output.
@@
The artifacts associated with the autoencoder-based CoNNear models were systematically examined across each auditory processing stage, which include CoNNearcochlear, CoNNearIHC and CoNNearANF stages as part of the closed-loop system described in [drakopoulos2023neural]. These artifacts are identified and quantified using the deterministic auditory computational models [verhulst2018computational], which served as targets during prior CoNNear training. Frequency responses for a 1-kHz tonal input across these stages are illustrated in Fig. 4(a-c). Fig. 4d shows the output of the HA model trained with the prior CoNNear framework to compensate for a high-frequency sloping hearing loss [drakopoulos2023neural]. The prior CoNNear-based models exhibit spectral peaks in the magnitude spectrum of the cochlear, IHC, and ANF responses that are not present in the target model. During the training of the HA model, which minimizes the difference between NH and HI AN responses, the HA models incorporated tonal artifacts as shown in Fig. 4d. These artifacts degrade the resulting audio quality and should be systematically excluded from the closed-loop framework.
@@
Figure 2: Generic diagram of the closed-loop framework for designing individualized hearing-aid algorithms. The three auditory modules—Cochlear, IHC, and ANF—are implemented using deep neural network architectures (e.g., CoNNear and dCoNNear).
@@ III Artifact-free dCoNNear-based closed-loop system
The artifact-free dCoNNear-based closed-loop framework for individualized DNN-based hearing-aid (HA) model training is illustrated in Fig. 2. This framework features two pathways: one for the AN response ($r_{f}$) of a normal hearing (NH) system, and the other for the response ($\hat{r_{f}}$) of a hearing-impaired (HI) system. Each pathway includes three primary components corresponding to distinct auditory processing stages, the cochlea, inner hair cells (IHC), and auditory nerve fibers (ANF). The NH AN response ($r_{f}$) is generated by combining three types of fibers (high-spontaneous rate (HSR), medium-spontaneous rate (MSR), and low-spontaneous rate (LSR)). These fibers are weighted by $H_{NH}$, $M_{NH}$, and $L_{NH}$, reflecting the typical innervation of each IHC in a healthy auditory system. To simulate hearing-impaired profiles, we adjust these weights to $H_{HI}$, $M_{HI}$, and $L_{HI}$ to reflect the effects of cochlear synaptopathy (CS). The resulting AN responses ($r_{f}$) and ($\hat{r_{f}}$) provide biophysically realistic time-frequency representations of sound (neurograms), simulated at different cochlear locations. These neurograms reflect instantaneous firing rates across cochlear channels with center frequencies (CFs) ranging from 112 Hz to 12 kHz [greenwood1990cochlear]. To tailor the hearing-impaired periphery to an individual’s sensorineural hearing loss (SNHL) profile, adjustments are made using audiometry to simulate the outer-hair-cell (OHC) damage in the cochlea or using auditory evoked potentials [keshishzadeh2021towards] to estimate CS. This involves introducing frequency-dependent OHC loss in CoNNearcochlear and/or CS in the CoNNearANF model. The DNN-based HA models are subsequently trained by minimizing a predefined loss function between the simulated NH and HI responses [drakopoulos2023neural]. All auditory elements and HA models within this framework are built upon the dCoNNear topology.
@@
To overcome the limitations of conventional encoder-decoder architectures, we introduce the dCoNNear model (Fig. 5), which explicitly addresses structural artifacts, such as aliasing, tonal distortions, and imaging effects, commonly introduced by suboptimal downsampling and upsampling methods [pons2021upsampling, zhang2019making, shang2023analysis]. To eliminate the architectural causes of these artifacts, dCoNNear replaces the encoder-decoder structure with a fully convolutional stack of FIR-like memory blocks as illustrated in Fig. 5 (b). Each block employs dilated depthwise convolutions to capture long-term temporal dependencies without altering the temporal resolution. By avoiding any explicit downsampling or upsampling operations, the architecture removes the structural conditions under which aliasing and spectral folding arise, thereby preventing the emergence of these artifacts in the output signal. This architecture maintains the ability to capture long-range dependencies while substantially reducing artifacts.
The overall architecture, shown in Fig. 5(a), consists of $M$ memory blocks with dilation factors $2^{m-1}$, repeated R times, resulting in a total of N = M * R blocks. Residual connections between memory blocks facilitate learning by addressing the vanishing gradient problem. Skip connections aggregate the outputs from each memory block through weighted summation, and these aggregated outputs are then processed by a non-linear activation function followed by a convolutional layer to produce the final model outputs [conv-tasnet].
Each memory block (Fig. 5(b)) begins with a pointwise convolution, followed by two dilated depthwise convolutions—Dh-conv and Df-conv—which capture historical and future temporal context, respectively. Depthwise convolution applies a single filter per input channel, reducing the number of parameters and helping capture long-term dependencies while maintaining a compact model size. The outputs from the pointwise and depthwise convolution layers are combined, passed through an activation function, and directed to subsequent blocks and skip connection paths. The learnable FIR-like memory block [zhang2018DFSMN] is formulated as: